1. Convert the following infix notations to postfix notations. (15%)

(1) A×B+C÷D

(2) A+B+C+D

(3) A+B×(C-D)

(4) (A+B+C)×D

(5) A+B×C+D
Answer #1

(1)

Given A×B+C÷D

=((A×B)+(C÷D))

=(ABx)+(CD÷)

=ABxCD÷+

(2)

Given A+B+C+D

=(((A+B)+C)+D)

=((AB++C)+D)

=AB+C++D

=AB+C+D+

(3)

Given A+B×(C-D)

=A+(Bx(C-D))

=A+(Bx(CD-))

=A+(BCD-x)

=ABCD-x+


(4) Given (A+B+C)×D

=(((A+B)+C)×D)

=((A+B)+C)Dx

=((AB+)+C)Dx

=AB+C+Dx

(5) Given A+B×C+D

= ((A+(B×C))+D)

=((A+BCx)+D)

=(ABCx+)+D)

=ABCx+D+